# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v2

@@ 3.3 Initialization of Reborn
In Stage 1, the segmentation model depends on the phoneme prediction when calculating the rewards. As a result, Reborn cannot work without properly initializing the phoneme prediction model. We use wav2vec-U to train a phoneme prediction model and use it as the initialization for the phoneme prediction model in Reborn. We briefly introduce wav2vec-U in Appendix D.
@@ 4.2 Evaluation Metrics
We use phoneme error rate (PER) and word error rate (WER) to evaluate the ASR performance. If not specified, we use greedy decoding to obtain phoneme-level results. For decoding word-level outputs, we perform WFST decoding [36, 37] using PyKaldi [8]. We leave the details in Appendix E.7.
@@ 4.3 Implementation Details
For the English datasets, we use wav2vec 2.0 [3] [^1] https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_new.pt to extract speech features from speech signals; for MLS, we use XLSR-53 [13] [^2] https://dl.fbaipublicfiles.com/fairseq/wav2vec/xlsr_53_56k.pt as the feature extractor. Both of them can be found in fairseq [41]. The 4-gram phoneme LM is derived using KenLM [21]. We describe more details about the data preparation and LM formulation in Appendix E.1, which basically follows wav2vec-U [4]. The segmentation model is a two-layer one-dimensional CNN model, detailed in Appendix E.3. For LibriSpeech and TIMIT, we train Reborn for two iterations. For MLS, we only train one iteration since we do not find the performance to improve in the second iteration.
@@ 5.1 Main Results
Table 1: PER/WER on LibriSpeech using 100 hours speech data. $\dagger$: Our reproduction. (wav2vec-U and wav2vec-U 2.0 only report results of using 960 hours of unlabeled speech). HMM ST indicates HMM self-training (Appendix E.2).
@@
Reborn significantly outperforms all prior methods on LibriSpeech
Our experiment on LibriSpeech follows Gao et al. [17] to use the 100-hour training split of LibriSpeech as the unlabeled speech data. Compared with all the baselines in Table 1, Reborn achieves the lowest PER and WER. wav2vec-U and EURO both use hand-crafted rules to obtain the segmental boundaries, while wav2vec-U 2.0 removes the feature segmentation steps. The superior performance of Reborn illustrates the importance of learning the segmental boundaries tailored for the phoneme prediction model. Notably, without using HMM self-training, Reborn already has PER/WER lower than wav2vec-U 2.0 with HMM self-training, which is the prior SoTA method on LibriSpeech. In Appendix C.5, we further apply the Reborn pipeline with speech foundation models other than wav2vec, and we find that Reborn yields notable performance improvement when using HuBERT [22] or WavLM [11] as the speech foundation model, illustrating the generalizability of the pipeline.
Self-training further improves PER/WER of Reborn
Self-training uses the phoneme predictor’s prediction as the pseudo-label and trains a new phoneme prediction model using the pseudo-label as the training data. It is commonly used in UASR to boost the performance [4, 28]. In Table 1, we show that Reborn can be integrated with Hidden Markov Models (HMM) self-training (Appendix E.2) to further lower the PER/WER. We reach a new SoTA on LibriSpeech under the setting of 100-hour speech data, outperforming the prior SoTA in PER and WER by 5% and 6%, respectively. Surprisingly, Reborn with HMM self-training outperforms training wav2vec-U with the oracle boundary. This shows that Reborn can be combined with self-training to improve performance effectively. Due to limited computation resources, we only conduct self-training on LibriSpeech.
Table 3: PER/WER on MLS. $\dagger$: Results from Baevski et al. [4]. *: Our reproduction of wav2vec-U, used as the initialization of the phoneme prediction model in Reborn. $\ddagger$: Results from Liu et al. [28].
@@
Boundary method Prec. Rec. F1 Freq. PER (%) (a) Oracle 1.0 1.0 1.0 11.89 13.8 (b) $k$-means-based 0.64 0.77 0.70 14.27 19.3 (c) Strgar and Harwath [53] 0.75 0.75 0.75 12.26 23.2 (d) Reborn 0.57 0.78 0.65 16.29 12.9
Reborn outperforms all prior UASR methods on TIMIT
Table 2 shows the PER of TIMIT, and we again find that Reborn achieves the lowest PER compared with all prior UASR methods. This indicates that Reborn not only works on large datasets like LibriSpeech but also works on small datasets like TIMIT, which only contains about 3 hours of audio data for training. Even using greedy decoding only, Reborn outperforms the prior best-performing UASR model (row (e)), which relies on prefix decoding. EURO shows that replacing the feature extractor with WavLM [11] (row (e)) outperforms using wav2vec 2.0 (row (d)) as the speech feature extractor. Since Reborn using wav2vec 2.0 already outperforms EURO with WavLM, we leave changing the feature extractor in Reborn on TIMIT as future work.
Reborn performs well on languages other than English
The results on MLS are presented in Table 3. We find out that a single iteration is enough for performance convergence on MLS. Recall that Reborn is initialized from wav2vec-U, but Reborn outperforms wav2vec-U by a large margin in all six languages. This large performance gap underlines the importance of learning the boundaries of the segmental boundaries. Moreover, Reborn outperforms wav2vec-U 2.0 except for Portuguese, and the average WER on the six languages is 3.9% lower than wav2vec-U 2.0.
@@ 5.2 Boundary Analysis
The core of Reborn is the segmentation model that learns the segmental structure’s boundaries. In this section, we look at how the quality of segmental boundaries affects the performance of phoneme prediction. We compare the boundaries obtained by five methods: (a) the oracle phoneme boundaries obtained by forced alignment [34], which requires paired speech-text data to learn the alignment; (b) the $k$-means-based segmentation boundary in wav2vec-U; (c) the phoneme boundary obtained by the SoTA unsupervised phoneme segmentation method [53] on LibriSpeech; (d) the boundaries learned by the segmentation model of Reborn in the first iteration before boundary merging.
First, we focus on the PER in Table 4 when using different segmentation methods. The PER is obtained by taking the four different segmented features to the identical phoneme prediction model: the phoneme prediction model trained using the $k$-means-based segmentation of wav2vec-U. We find that simply replacing the $k$-means-based boundary with the oracle phoneme boundary reduces the PER by 5.5%, showing the imperfect hand-crafted $k$-means-based segmentation is the bottleneck of UASR. Next, even when using the boundary predicted by the SoTA unsupervised phoneme segmentation [53], the PER does not reduce. Conversely, the boundary learned by Reborn achieves the lowest PER. This is because the boundary learned by Reborn is directly tailored for the phoneme prediction model.
Next, we discuss the phoneme boundary quality in Table 4. The phoneme boundary quality is evaluated with boundary precision, recall, and F1 with a 20ms tolerance window, following the harsh scheme from Strgar and Harwath [53]. We leave the discussion about the different schemes for boundary evaluation in Appendix G. We also report the average number of segments per second (Freq. in Table 4). Interestingly, the boundary F1s of the oracle boundary (row (a)) and Strgar and Harwath [53] (row (c)) are much higher than the boundary F1 of Reborn (row (d)), but the PER of Reborn is much better. This is because Reborn’s segmentation model predicts more segments than the number of phonemes in the utterance (before boundary merging), indicating that it learns some segmental structures smaller than the phonemes. This can be seen from the lower precision (0.57) and higher frequency (16.29) compared to oracle boundaries. However, segmenting the speech feature into smaller units is not problematic because even if the consecutive speech features of the same phonemes are split into two segments by the segmentation model, as long as they have the same phoneme prediction, the duplicated phoneme prediction will be removed in the de-duplication step, and thus not affect the PER.
@@
Figure 2: PER of each stage during Reborn’s two-stage iterative training on the test-clean split of LibriSpeech. St.: stage; w2vu: wav2vec-U
@@ 5.3.1 Reward Design and Behavior Cloning
We verify the effectiveness of our three proposed reward functions and behavior cloning initialization in Table 5. While the perplexity difference reward alone is sufficient to guide the segmentation model, adding edit distance and length difference rewards improves performance further. As previously mentioned, using only the perplexity difference reward will make the segmentation model predict segments overly focusing on lowering PPL. And a phoneme sequence with a lower perplexity may not always be a better transcription. To deal with such overfitting problem, we design the two regularization rewards, namely, $R_{edit}$ and $R_{len}$, to make the model learn to segment for lower perplexity while grounded on a reasonable transcription from the previous iteration. Our results in Table 5 further evidence the effectiveness of the two rewards. Although having a higher perplexity, using the additional regularized rewards has successfully increased the performance with lower PER (row (c) vs. row (a)). We also find that removing BC (row (d)) worsens PER compared with using BC (row (c)). Additionally, BC greatly speeds up RL training convergence, as shown in Figure 3.
@@ 5.3.2 Iterative Training
In Figure 2, we show the PER on LibriSpeech test-clean split during the iterative training process of Reborn. We observe that after training the segmentation model in the first iteration, the PER drastically drops by 6.4% compared with wav2vec-U, which is used to initialize the phoneme prediction model in iteration one. This shows that the quality of the segmental structure’s boundaries is important to the performance of UASR. Afterward, the PER of Reborn decreases steadily across iterative training, showing that training the segmentation model and the phoneme prediction model is critical to improving the performance of Reborn. We find that the PER does not drop significantly after the fifth iteration. Additionally, we provide more evidence to show that the phoneme predictors are iteratively polished in Appendix C.4.
6 Conclusion
@@
In unsupervised ASR, learning the mapping between the segmental structures in speech and their corresponding transcription is challenging, especially without paired data. To tackle this, we propose Reborn, an iterative training method for unsupervised ASR. Reborn iteratively trains the segmentation model and the phoneme prediction model to improve the ASR performance. The segmentation model is trained using RL with carefully designed rewards to guide the segmentation model to find the boundaries of the segmental structures in speech signals. Extensive experiments on three datasets spanning seven languages show that Reborn outperforms prior best-performing unsupervised ASR models in all datasets except one language in MLS. We conduct comprehensive ablation studies to show that each component in Reborn is critical to the performance. We also explain the effectiveness of Reborn by analyzing its segmentation patterns and find that Reborn tends to produce segmental structures smaller than phonemes, which helps the generators to better predict the phoneme transcription.
Appendix / supplemental material
Appendix A Limitations
@@
Recent advancements in unsupervised ASR are still in the developmental stages, and we have not yet implemented Reborn in more realistic or noisy environments. Additionally, the iterative training paradigm may amplify any existing biases in the dataset, an aspect that remains unexplored in this study. Furthermore, if the phoneme prediction model is poorly initialized, Reborn may not be able to provide huge performance improvements. In this work, lexicons are required to perform phonemicization on the unpaired text. Consequently, learning the unsupervised ASR system directly from speech to word-level tokens remains an open problem.
Appendix B Broader Impacts
@@
Unsupervised automatic speech recognition (UASR) - predicting the textual transcriptions for the given speech signals using only unpaired speech and text data - is highly attractive. The existence of thousands of low-resourced languages spoken globally is the major reason. This paper presents Reborn, a novel unsupervised ASR training algorithm. We foresee that Reborn has great potential to make low-resource languages more accessible to train ASR systems, making speech technology more accessible and boundaryless.
Appendix C Additional Results
C.1 Impact of Behavior Cloning on Model Convergence
Figure 3: Comparison of training dynamics between BC pre-trained and from-scratch models. This figure shows the evolution of Phoneme Error Rate (PER) across training epochs on the test-clean split of LibriSpeech, illustrating the faster convergence and improved efficiency achieved with BC pre-training.
In this ablation study, we investigate the impact of Behavior Cloning (BC) pre-training on the performance of our segmenter model. Our results show that BC pre-training enhances sample efficiency and accelerates convergence compared to training from scratch. The training curves in Figure 3 highlight these differences, showing both faster convergence and improved performance with BC pre-training.
The segmenter solely pre-trained through BC exhibits noticeable improvements, but this is not yet substantial in terms of Phoneme Error Rate (PER) initially. However, this pre-training boosts the model’s boundary prediction capabilities, accelerating learning in later reinforcement learning phases. This indicates that BC pre-training is essential for enhancing our unsupervised automatic speech recognition framework, making it a crucial optimization step.
C.2 Reward Design Details

### v3

@@ 3.3 Initialization of Reborn
In Stage 1, the segmentation model depends on the phoneme prediction when calculating the rewards. As a result, Reborn cannot work without properly initializing the phoneme prediction model. We use wav2vec-U to train a phoneme prediction model and use it as the initialization for the phoneme prediction model in Reborn. We briefly introduce wav2vec-U in Appendix B.
@@ 4.2 Evaluation Metrics
We use phoneme error rate (PER) and word error rate (WER) to evaluate the ASR performance. If not specified, we use greedy decoding to obtain phoneme-level results. For decoding word-level outputs, we perform WFST decoding [38, 39] using PyKaldi [9]. We leave the details in Appendix C.7.
@@ 4.3 Implementation Details
For the English datasets, we use wav2vec 2.0 [3] to extract speech features from speech signals; for MLS, we use XLSR-53 [14] as the feature extractor. Both of them can be found in fairseq [44]. The 4-gram phoneme LM is derived using KenLM [22]. We describe more details about the data preparation and LM formulation in Appendix C.1, which basically follows wav2vec-U [4]. All the trainable models, including the segmentation model, the phoneme prediction model, and the discriminator in GAN training, are composed of one-dimensional CNN layers, as detailed in Appendix C.3. For LibriSpeech and TIMIT, we train Reborn for two iterations. For MLS, we only train one iteration since we do not find the performance to improve in the second iteration.
@@ 5.1 Main Results
Table 1: PER/WER on LibriSpeech using 100 hours speech data. $\dagger$: Our reproduction. (wav2vec-U and wav2vec-U 2.0 only report results of using 960 hours of unlabeled speech). HMM ST indicates HMM self-training (Appendix C.2).
@@
Reborn significantly outperforms all prior methods on LibriSpeech. Our experiment on LibriSpeech follows Gao et al. [18] to use the 100-hour training split of LibriSpeech as the unlabeled speech data. Compared with all the baselines in Table 1, Reborn achieves the lowest PER and WER. wav2vec-U and EURO both use hand-crafted rules to obtain the segmental boundaries, while wav2vec-U 2.0 removes the feature segmentation steps. The superior performance of Reborn illustrates the importance of learning the segmental boundaries tailored for the phoneme prediction model. Notably, without using HMM self-training, Reborn already has PER/WER lower than wav2vec-U 2.0 with HMM self-training, which is the prior state-of-the-art (SoTA) method on LibriSpeech. In Appendix A.4, we further apply the Reborn pipeline with speech foundation models other than wav2vec 2.0. We find that Reborn yields notable performance improvement when using HuBERT [23] or WavLM [12] as the feature extractor, illustrating the generalizability of our method.
Self-training further improves PER/WER of Reborn. Self-training uses the phoneme predictor’s prediction as the pseudo-label and trains a new phoneme prediction model using the pseudo-label as the training data. It is commonly used in UASR to boost the performance [4, 30]. In Table 1, we show that Reborn can be integrated with Hidden Markov Models (HMM) self-training (Appendix C.2) to further lower the PER/WER. We reach a new SoTA on LibriSpeech under the setting of 100-hour speech data, outperforming the prior SoTA in PER and WER by 5% and 6%, respectively. Surprisingly, Reborn with HMM self-training outperforms training wav2vec-U with the oracle boundary. This shows that Reborn can be combined with self-training to improve performance effectively. Due to limited computation resources, we only conduct self-training on LibriSpeech.
Table 3: WER on MLS. $\dagger$: Results from Baevski et al. [4]. *: Our reproduction of wav2vec-U, used as the initialization of the phoneme prediction model in Reborn. $\ddagger$: Results from Liu et al. [30].
@@
Boundary method Prec. Rec. F1 Freq. PER(a) Oracle 1.0 1.0 1.0 11.89 13.8 (b) $k$-means-based 0.64 0.77 0.70 14.27 19.3 (c) Strgar and Harwath [56] 0.75 0.75 0.75 12.26 23.2 (d) Reborn 0.57 0.78 0.65 16.29 12.9
Reborn outperforms all prior UASR methods on TIMIT. Table 2 shows the PER of TIMIT, and we again find that Reborn achieves the lowest PER compared with all prior UASR methods. This indicates that Reborn not only works on large datasets like LibriSpeech but also works on small datasets like TIMIT, which only contains about 3 hours of audio data for training. Even using greedy decoding only, Reborn outperforms the prior best-performing UASR model (row (e)), which relies on prefix decoding. EURO shows that replacing the feature extractor with WavLM [12] (row (e)) outperforms using wav2vec 2.0 (row (d)) as the speech feature extractor. Since Reborn using wav2vec 2.0 already outperforms EURO with WavLM, we leave changing the feature extractor in Reborn on TIMIT as future work.
Reborn performs well on languages other than English. The results on MLS are presented in Table 3. We find out that a single iteration is enough for performance convergence on MLS. Recall that Reborn is initialized from wav2vec-U, but Reborn outperforms wav2vec-U by a large margin in all six languages. This large performance gap underlines the importance of learning the boundaries of the segmental boundaries. Moreover, Reborn outperforms wav2vec-U 2.0 except for Portuguese, and the average WER on the six languages is 3.9% lower than wav2vec-U 2.0.
@@ 5.2 Boundary Analysis
The core of Reborn is the segmentation model that learns the segmental structure’s boundaries. In this section, we look at how different segmental boundaries affect the performance of phoneme prediction. We compare the boundaries obtained by four methods: (a) the oracle phoneme boundaries obtained by forced alignment [36], which requires paired speech-text data to learn the alignment; (b) the $k$-means-based segmentation boundary in wav2vec-U; (c) the phoneme boundary obtained by the SoTA unsupervised phoneme segmentation method [56] on LibriSpeech; (d) the boundaries learned by the segmentation model of Reborn in the first iteration before boundary merging.
First, we focus on the PER in Table 4 when using different segmentation methods. The PER is obtained by taking the four different segmented features to the identical phoneme prediction model: the phoneme prediction model trained using the $k$-means-based segmentation of wav2vec-U. Note that the phoneme prediction model here is unsupervised and non-ideal. We find that simply replacing the $k$-means-based boundary with the oracle phoneme boundary reduces the PER by 5.5%, showing the imperfect hand-crafted $k$-means-based segmentation is the bottleneck of UASR. Next, even when using the boundary predicted by the SoTA unsupervised phoneme segmentation [56], the PER does not reduce. Conversely, the boundary learned by Reborn achieves the lowest PER. This is because the boundary learned by Reborn is directly tailored for the phoneme prediction model.
Next, we discuss the phoneme boundary results in Table 4. The phoneme boundary results is evaluated with boundary precision, recall, and F1 with a 20ms tolerance window, following the harsh scheme from Strgar and Harwath [56]. We leave the discussion about the different schemes for boundary evaluation in Appendix E. We also report the average number of segments per second (Freq. in Table 4). Interestingly, the boundary F1s of the oracle boundary (row (a)) and Strgar and Harwath [56] (row (c)) are much higher than the boundary F1 of Reborn (row (d)), but the PER of Reborn is much better. This is because Reborn’s segmentation model predicts more segments than the number of phonemes in the utterance (before boundary merging), indicating that it learns some segmental structures smaller than the phones. This can be seen from the lower precision (0.57) and higher frequency (16.29) compared to oracle boundaries. However, segmenting the speech feature into smaller units is not problematic because even if the consecutive speech features of the same phonemes are split into two segments by the segmentation model, as long as they have the same phoneme prediction, the duplicated phoneme prediction will be removed in the de-duplication step, and thus not affect the PER.
@@
Figure 2: PER across training epochs on the test-clean split of LibriSpeech. BC pretraining speeds up convergence and raises performance.
@@ 5.3.1 Reward Design and Behavior Cloning
We verify the effectiveness of our three proposed reward functions and behavior cloning initialization in Table 5. While the perplexity difference reward alone is sufficient to guide the segmentation model, adding edit distance and length difference rewards improves performance further. As previously mentioned, using only the perplexity difference reward will make the segmentation model predict segments overly focusing on lowering PPL. And a phoneme sequence with a lower perplexity may not always be a better transcription. To deal with such overfitting problem, we design the two regularization rewards, namely, $R_{\text{edit}}$ and $R_{\text{len}}$, to make the model learn to segment for lower perplexity while grounded on a reasonable transcription from the previous iteration. Our results in Table 5 further evidence the effectiveness of the two rewards. Additionally, we observe that removing BC leads to a decline in PER compared to using BC (see row (d) vs. row (c)). To deepen this analysis, we present the PER curve on the LibriSpeech test-clean set with and without BC. The results in Figure 2 indicate that BC helps enhance performance and accelerate convergence.
@@ 5.3.2 Iterative Training
In Figure 3, we show the PER on LibriSpeech test-clean split during the iterative training process of Reborn. We observe that after training the segmentation model in the first iteration, the PER drastically drops by 6.4% compared with wav2vec-U, which is used to initialize the phoneme prediction model in iteration one. This shows that the quality of the segmental structure’s boundaries is important to the performance of UASR. Afterward, the PER of Reborn decreases steadily across iterative training, showing that training the segmentation model and the phoneme prediction model is critical to improving the performance of Reborn. We find that the PER does not drop significantly after the fifth iteration. Last but not least, we provide more evidence in Appendix A.3 to showcase that the phoneme predictors are iteratively polished.
Figure 3: PER of each stage during Reborn’s two-stage iterative training on the test-clean split of LibriSpeech. St.: stage; w2vu: wav2vec-U.
Table 6: The boundary evaluation results on TIMIT. Reborn achieves better boundary evaluation scores than the original $k$-means-based method, and is comparable with Strgar and Harwath [56]’s after boundary merging. *: from literature. All the metrics are calculated on the test-all split.
[…]
5.3.3 Boundary Evaluation on TIMIT
In Section 5.2, we demonstrate that Reborn segmentation delivers fine-grained segments, achieving the highest performance gain among segmentation methods given the same non-ideal phoneme prediction model. Here, we extend our boundary evaluation results to TIMIT, a smaller dataset with human-annotated phone boundaries, to provide more comprehensive boundary insights. Table 6 shows that the initial boundaries learned by Reborn, optimized explicitly for the phoneme prediction model, already achieve a high recall. With boundary merging, i.e., consecutive segments with the same phoneme prediction are merged as illustrated in Figure 1-(c), Reborn achieves results close to  Strgar and Harwath [56]’s. In line with the previous context, we emphasize that Reborn achieves substantial performance improvement in PER by tailoring segmentation to the non-ideal unsupervised phoneme prediction model, rather than solely aiming for higher boundary evaluation scores. This represents a critical and intriguing finding in our work.
6 Discussion
@@
In unsupervised ASR, learning the mapping between the segmental structures in speech and their corresponding transcription is challenging, especially without paired data. To tackle this, we propose Reborn, an iterative training method for unsupervised ASR. Reborn iteratively trains the segmentation model and the phoneme prediction model to improve the ASR performance. The segmentation model is trained using RL with carefully designed rewards to guide the segmentation model to find the boundaries of the segmental structures in speech signals. Extensive experiments on three datasets spanning seven languages show that Reborn outperforms prior best-performing unsupervised ASR models in all datasets except one language in MLS. We conduct comprehensive ablation studies to show that each component in Reborn is critical to the performance. We also explain the effectiveness of Reborn by analyzing its segmentation patterns and find that Reborn tends to produce segmental structures smaller than phones, which helps the generators predict the phoneme transcription better.
@@
Limitations. Recent advancements in unsupervised ASR are still in the developmental stages, and we have not yet implemented Reborn in more realistic scenarios, such as low-resource languages or noisy environments. Additionally, the iterative training paradigm may amplify any existing biases in the dataset, an aspect that remains unexplored in this study. Furthermore, if the phoneme prediction model is poorly initialized, Reborn may not be able to provide huge performance improvements. In this work, lexicons are required to perform phonemicization on the unpaired text. Consequently, learning the unsupervised ASR system directly from speech to word-level tokens remains an open problem, where recent works aim to tackle [30, 41].
@@
Broader impacts. Unsupervised automatic speech recognition (UASR) - predicting the textual transcriptions for the given speech signals using only unpaired speech and text data - is highly attractive. The existence of thousands of low-resourced languages spoken globally is the major reason. This paper presents Reborn, a novel unsupervised ASR training algorithm. We foresee that Reborn has great potential to make low-resource languages more accessible to train ASR systems, making speech technology more accessible and boundaryless.
Acknowledgment
We thank the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources. Shao-Hua Sun was supported by the Yushan Fellow Program by the Ministry of Education, Taiwan. We also extend our gratitude to Alexei Baevski, Wei-Ning Hsu, and Alexander H. Liu, authors of wav2vec-U and wav2vec-U 2.0, for providing valuable insights into their work.
Appendix
Appendix A Additional Results
Table 7: The table for stability analysis. Each method is trained for 5 runs on LibriSpeech and evaluated on the test-clean split. Reborn performs steadily well while being fully converged in both of the stages. The ablations show the effectiveness of applying boundary merging and parameter initialization from the previous iteration for the stage 2 GAN training. The abbreviation "adj." indicates adjacent pooling, which is the 2nd stage of pooling in wav2vec-U (see Appendix B). The "Freq." in the last column is the number of boundaries per second.
A.1 Stability Analysis
Next, we focus on the ablations of Reborn stage 2 GAN training in Table 7. When the boundary merging is not applied, the GAN training on the original Reborn-segmented features is still better than the initial stage. However, the average PER and the standard deviation are larger. The observation indicates that the training is slightly unstable, but does not diverge too much since it is still fully converged. On the other hand, when we train the next-iteration GAN with completely random initialized parameters, the performance degrades even if we train the model for more steps. These two ablations directly demonstrate the effectiveness of adopting boundary merging and parameter initialization from the previous iteration, while the GAN training can still work on the original boundary or with randomly initialized parameters.
A.2 Reward Design Details